40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 9 17:12:52 2020, gautam, Update, BHD, In-air BHD OMCchamber.pdfAP_Table_20180328.pdf
    Reply  Tue Jul 14 15:29:25 2020, gautam, Update, BHD, In-air BHD - DCPD amplifier noise DCPDamp.pdf
    Reply  Wed Jul 15 19:11:40 2020, gautam, Update, BHD, In-air BHD - alignment into OMC 
       Reply  Thu Jul 16 01:12:22 2020, gautam, Update, BHD, In-air BHD - preparing the LO path IMG_8626.JPGIMG_8627.JPGIMG_8628.JPGIMG_8629.JPG
          Reply  Sun Jul 19 15:40:15 2020, gautam, Update, BHD, In-air BHD - CDS and wiring summary wiringDiagram.pdf
             Reply  Wed Jul 29 11:57:59 2020, rana, Update, BHD, In-air BHD - CDS and wiring summary 
          Reply  Mon Jul 20 17:55:15 2020, gautam, Update, BHD, In-air BHD - preparing the LO path 
             Reply  Tue Jul 21 00:30:24 2020, gautam, Update, BHD, In-air BHD - LO RIN LO_RIN.pdf
Message ID: 15495     Entry time: Mon Jul 20 17:55:15 2020     In reply to: 15489     Reply to this: 15497
Author: gautam 
Type: Update 
Category: BHD 
Subject: In-air BHD - preparing the LO path 

The LO pickoff has been coupled into a fiber with ~90% MM (8 mW / 9 mW input). While I wait for the DCPD electronics to be found in the Cryo lab, I want to monitor the stability of the pointing, polarization etc, so I'd like to clear some space on the AP table that was occupied for the mode spectroscopy project. If there are no objections before 2pm tomorrow July 21 2020, I will commence this work.

ELOG V3.1.3-